OpenStudy (anonymous):

What is the value of x in the following system: y=2x+5 y=-3x+10

OpenStudy (anonymous):

ok.... since both equations start off as "y=..." set them equal to each other....

OpenStudy (anonymous):

what do you mean equal to each other?

OpenStudy (anonymous):

2x+5 = -3x+10

OpenStudy (anonymous):

OHHHHH!! OK I got x=-3

OpenStudy (anonymous):

no... try again.... be careful...

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

how did u get that? try again...

OpenStudy (anonymous):

okay so i got 2x+15=-3x and then i subtracter -2 from -3 on the other side

OpenStudy (anonymous):

ok... you moved the 10 over from the right side so that means you needed to SUBTRACT 10 from both sides... that 15 should be a -5... ????

OpenStudy (anonymous):

OHHHHH! SORRY!! I forgot to do inverse operation !

OpenStudy (anonymous):

so let's take it from there now.... 2x - 5 = -3x what next?

OpenStudy (anonymous):

you subtract -2 from -3x

OpenStudy (anonymous):

you mean subtract 2x from both sides.... ??? 2x - 5 -2x = -3x -2x

OpenStudy (anonymous):

can you simplify both sides??

OpenStudy (anonymous):